OpenStudy (anonymous):

Nancy is a professional skydiver. For her next trip, she is planning to jump from a plane at 13,000 feet above sea level. She will be landing in a valley located at 120 feet below sea level. If Nancy opens her shoot after falling 10,500 feet from the plane, how much farther will she fall, in feet, to reach the valley floor?

OpenStudy (hanner_b_nanner):

She will have 2,620 more feet to go. 13,000+120=13,120-10,500=2,620. :)